oxidation state Aluminum in AlCl3, Carbon in C2O4^2-,oxygen in O2
Answer #1

тя. Oidation state = hypothetical charge on atoms. * AlCl3 x + 3x(-1) =0 تر X x -3 = 0 =+3 Al +3 = 2042- 2x2 + 4x(-2) = -2 2